The world's most populous countries are China and India. As of mid-2005, the combined population of these two countries was estimated at 2.4 billion. If there was about 4/5 as many people living in India as China, what was the population of each country, to the nearest tenth of a billion?
Answer by stanbon(75887)

The world's most populous countries are China and India. As of mid-2005, the combined population of these two countries was estimated at 2.4 billion. If there was about 4/5 as many people living in India as China, what was the population of each country, to the nearest tenth of a billion?
---------------
I + C = 2.4 billion
I= (4/5)C
----------
Substitute and solve for C
(4/5)C + C = 2.4 billion
1.8C = 2.4
C = 1 1/3 is appox 1.3 billion
-------------------
I = (4/5)(4/3) = 16/15 = 1.0667 is approx 1.1 billion
=======================
Cheers,
Stan H.